OpenStudy (anonymous):

What is the simplified form of (4x^2-25)/(2x+5) 2x-5, with the restriction x≠ -5/2 <--? 2x+5, with the restriction x≠ 5/2 2x-5, with the restriction x≠ 5/2 2x+5, with the restriction x≠ -5/2

OpenStudy (freckles):

first of all we cannot divide by 0 so when is 2x+5, 0? that is solve 2x+5=0

OpenStudy (freckles):

this will give you your restriction

OpenStudy (freckles):

then you will need to simplify your expression hint difference of squares a^2-b^2=(a-b)(a+b)